# Characterization of Residual Charge Images in LSST Camera e2v CCDs

### v1

@@ 5 CONCLUSION
LSSTCam e2v CCDs exhibit residual charge effects in images following exposure to bright sources. These manifest not just as ghosts in the same region of the focal plane as the source, but also as a trail between the source and the serial register. This indicates charge is being captured in the CCD and likely on the CCD’s oxide surface layer interface not just during integration, but also during the parallel transfer and readout. Both these regions, but especially the trail residual charge images will create a large systematic error in the Rubin Observatory Legacy Survey of Space and Time if not addressed. The trail effect in particular would inject linear contrasting features on the scale of 10 electrons above the sky background which would greatly impact source extraction and photometry.
@@ Acknowledgments
We acknowledge significant help from Gregg Thayer, Yousuke Utsumi, Tony Johnson, Stuart Marshall and Pierre Antilogus. We thank Sean MacBride, Yousuke Utsumi and Andy Rasmussen for carefully reviewing our manuscript. This research is supported by DOE grant DE-SC0009999 and NSF grant AST-2205095.

### v2

@@ 5 CONCLUSION
LSSTCam e2v CCDs exhibit residual charge effects in images following exposure to bright sources. These manifest not just as ghosts in the same region of the focal plane as the source, but also as a trail between the source and the serial register. This indicates charge is being captured in the CCD and likely on the CCD’s oxide surface layer interface not only during integration, but also during the parallel transfer and readout. Both these regions, but especially the trail residual charge images will create a large systematic error in the Rubin Observatory Legacy Survey of Space and Time if not addressed. The trail effect in particular would inject linear contrasting features on the scale of 10 electrons above the sky background which would greatly impact source extraction and photometry.
@@ Acknowledgments
We acknowledge significant help from Gregg Thayer, Yousuke Utsumi, Tony Johnson, Stuart Marshall and Pierre Antilogus. We thank Sean MacBride, Yousuke Utsumi and Andy Rasmussen for carefully reviewing our manuscript.
This material is based upon work supported in part by the National Science Foundation through Cooperative Agreement AST-1258333 and Cooperative Support Agreement AST-1202910 managed by the Association of Universities for Research in Astronomy (AURA), and the Department of Energy under Contract No. DE-AC02-76SF00515 with the SLAC National Accelerator Laboratory managed by Stanford University. Additional Rubin Observatory funding comes from private donations, grants to universities, and in-kind support from LSST-DA Institutional Members
This research is supported by DOE grant DE-SC0009999 and NSF grant AST-2205095.